### v1

@@ Abstract
Results. By applying a 20 km s-1 threshold on the peak-to-peak radial velocities above which we would consider the star to be binary, the resulting observed spectroscopic binary fraction for our BSG sample is 23 3. An independent analysis of line profile variability reveals 11 (plus 5 candidates) double-lined spectroscopic binaries and 32 (plus 41 candidates) single-lined spectroscopic binaries. Based on these results, we estimated the overall observed binary fraction in this sample to be 34 3, which is close to the computed intrinsic binary fraction of 40 4. In addition, we derived reliable orbital periods for 41 spectroscopic binaries and potential binary candidates, among which there are 17 eclipsing binaries, including 20 SB1 and SB2 systems with periods of less than 10 d. We reported a significant drop in the binary fraction of BSGs with spectral types later than B2 and effective temperatures less than 18 kK, which could indicate the end of the main sequence phase in this temperature regime. We found no metallicity dependence in the binary fraction of BSGs, compared to existing spectroscopic surveys of the Galaxy and Large Magellanic Cloud.
@@
On the main sequence (MS), the progenitors of BSGs are OB-type stars. By now, it has been well established that OB stars have high multiplicity rates (50-70, see Sana et al., 2012; Banyard et al., 2022; Offner et al., 2023). However, for many decades, it was widely believed that BSGs are preferably single objects (Humphreys, 1978; McEvoy et al., 2015; Simón-Díaz et al., 2024; de Burgos et al., 2024). Thus, we would naturally expect that a significant fraction of BSGs are products of binary interaction. One of the explanations could be that, in the BSG regime, there is a high fraction of merger products (Vanbeveren et al., 2013; Justham et al., 2014; Menon et al., 2024; Schneider et al., 2024). One of the most famous example is the progenitor star of SN 1987A, which was expected to be a red supergiant (RSG, see Walborn et al., 1987); however, it was later shown the progenitor is a BSG and a binary-merger scenario provides a natural explanation for both the BSG star and the rings of ejected material surrounding it (Podsiadlowski, 1992; Podsiadlowski et al., 2007; Menon & Heger, 2017). Another hint of predominance of binary interaction products in the BSG domain is a low multiplicity fraction of BSGs claimed in different multi-epoch spectroscopic surveys across different metallicity environments, namely, the estimated observed binary fraction of BSGs in the Galaxy is in the range of 20-25 (see Simón-Díaz et al., 2024) and in the Large Magellanic Cloud (LMC), it is approximately 23 6 (Dunstall et al., 2015).
@@
To address these problems, we study the multiplicity properties of massive BSGs ( ) in the Small Magellanic Cloud (SMC) within the new Binarity at LOw Metallicity (BLOeM, Shenar et al., 2024) survey. Apart from homogeneous and statistically significant numbers of observed massive stars, this survey gives the unique opportunity to study the evolution of BSGs at low metallicity (). The BLOeM survey contains multi-epoch spectroscopic observations of nearly 1000 stars in the SMC across different spectral types. The survey has been divided into different sub-samples that represent various evolutionary stages of massive stars across the HRD, namely: the O-type star domain (Sana et al., subm. in Nature Astronomy), the B-type dwarfs (spectral types B0-B3, Villaseñor et al., subm. in A&A), OeBe stars (Bodensteiner et al., 2025), the BSGs (present work), and the cool supergiants (spectral types later than B3 up to F-type supergiants, see Patrick et al., 2025).
@@ 4.1 Multiplicity of BSGs
The overall determined spectroscopic binary fraction of BSGs in the SMC is similar to those detected in the LMC and in the Galaxy. Dunstall et al. (2015) analyzed multiplicity properties of 408 B-type stars in the LMC and reported the observed binary fraction among non-supergiants to be 25 2% and 23 6% for supergiants. In the Galaxy, 56 OB supergiants have been analyzed by Simón-Díaz et al. (2024) and the observed binary fraction was claimed to be 25%. Such an independence of BSG binary fraction with respect to the metallicity is in agreement with the binary fraction of early B-type eclipsing systems in the Galaxy, LMC, and SMC (within the ¡ 20 d, see Moe & Di Stefano, 2013). As Moe et al. (2019) argued, the explanation for this could be that significant accretion rates in the massive star domain are so high that the massive protostellar disks will always become gravitationally unstable and fragment independent of the metallicity environment (which is not the case in the solar-type star domain). Furthermore, in the early B-type giants and dwarf BLOeMs’ sample, it has been found that binary fraction to some extent is dependent on metallicity (for details, see Villaseñor et al. subm. in A&A). In the next subsection, we discuss the binary fraction as a function of SpT and luminosity.
@@ 4.2 The BSGs landscape on the HRD
In addition, if we look at the distribution of spectroscopic binaries for a given spectral type (see the left panel in Fig. 10), we can notice that the B2 domain contains the largest number of spectroscopic binaries (especially ”lpv/SB1”) with respect to the rest of the spectral types in our BSG sample. In total, there are up to 50 of potential spectroscopic binaries (apart from the evolutionary effect), which can also be explained by their relatively low luminosity; thus, some of these targets could belong to the B-dwarf sample, where the observed multiplicity fraction is significant (50, see the accompanying BLOeM paper Villaseñor et al. subm. in A&A). Also, some of these least luminous BSGs could be Cephei pulsators, as these tend to be more commonly found among late-O and early-B spectral types (Burssens et al., 2020); thus, this could explain a given overpopulation of lpv/SB1 targets.
@@
Another reason for non-uniform distribution could be the selection bias of BSGs with the luminosity class of II, indeed, in this domain there is a significant overlap with the B-giants BLOeM sample (i.e., luminosity class of III, Villaseñor et al. subm. in A&A), which could bias the overall number of BSGs in a given SpT and LC. Nevertheless, there is a significant drop in binary fraction below 18 kK and even though the present physical parameters of BSGs are not well defined yet, we can argue that it is a real drop in the binary fraction of BSGs at a given effective temperature and spectral type. Moreover, RV variations of late B-type supergiants are significantly smaller (less than 10 km s-1 for SpT B5 and later; see Patrick et al., 2025) than the ones studied in the present work, which confirms the absence of short-period supergiants at SpT later than B3. Nevertheless, the binary fraction among B3 BSGs is 20%, which agrees with the observed spectroscopic binary fraction of more evolved late-B, A-, and F-type supergiants (20% for the B5-9s SpT domain and significantly smaller binary fraction for the A- and F-type supergiants). It could indicate that at this SpT domain (B2) and regime, we observe the end of MS for a studied sample of BSGs; actually, it will also be interesting to see the behavior of sin for the same targets. If there are no fast-rotating ( sin ¡ 100 km s-1) BSGs with a SpT later than B2, it can also support the idea of the MS boundary at this evolutionary phase (as it has been shown in Vink et al., 2010). Notably, the BLOeM survey has been designed in a way to observe as many as possible bright massive stars in the SMC, while also aiming to achieve a homogeneous sampling across the G-band from 10 to 16.5 mag. The overall sample completeness for the stars with (i.e. those born as B-type stars) is 20% with respect to the Gaia catalogue within the range of 13 – 15 mag in the G-band (see Fig. 2 in Shenar et al., 2024). Thus, the given sample of BSGs and the other sub-samples of the BLOeM survey do not have a significant sample selection bias in terms of the magnitude and colour of targets that would affect the interpretation of our results because of the incompleteness of the samples.

### v2

@@ Abstract
Results. By applying a 20 km s-1 threshold on the peak-to-peak radial velocities above which we would consider the star to be binary, the resulting observed spectroscopic binary fraction for our BSG sample is 23 3. An independent analysis of line profile variability reveals 11 (plus 5 candidates) double-lined spectroscopic binaries and 32 (plus 41 candidates) single-lined spectroscopic binaries. Based on these results, we estimated the overall observed binary fraction in this sample to be 34 3, which is close to the computed intrinsic binary fraction of 40 4. In addition, we derived reliable orbital periods for 41 spectroscopic binaries and potential binary candidates, among which there are 17 eclipsing binaries, including 20 SB1 and SB2 systems with periods of less than 10 days. We reported a significant drop in the binary fraction of BSGs with spectral types later than B2 and effective temperatures less than 18 kK, which could indicate the end of the main sequence phase in this temperature regime. We found no metallicity dependence in the binary fraction of BSGs, compared to existing spectroscopic surveys of the Galaxy and Large Magellanic Cloud.
@@
On the main sequence (MS), the progenitors of BSGs are OB-type stars. By now, it has been well established that OB stars have high multiplicity rates (50-70, see Sana et al., 2012; Banyard et al., 2022; Offner et al., 2023). However, for many decades, it was widely believed that BSGs are preferably single objects (Humphreys, 1978; McEvoy et al., 2015; Simón-Díaz et al., 2024; de Burgos et al., 2024). Thus, we would naturally expect that a significant fraction of BSGs are products of binary interaction. One of the explanations could be that, in the BSG regime, there is a high fraction of merger products (Vanbeveren et al., 2013; Justham et al., 2014; Menon et al., 2024; Schneider et al., 2024). One of the most famous example is the progenitor star of SN 1987A, which was expected to be a red supergiant (RSG, see Walborn et al., 1987); however, it was later shown the progenitor is a BSG and a binary-merger scenario provides a natural explanation for both the BSG star and the rings of ejected material surrounding it (Podsiadlowski, 1992; Podsiadlowski et al., 2007; Menon & Heger, 2017). Another hint of predominance of binary interaction products in the BSG domain is a low multiplicity fraction of BSGs claimed in different multi-epoch spectroscopic surveys across different metallicity environments, namely, the estimated observed binary fraction of BSGs in the Galaxy is 27 (see, de Burgos et al., 2025) and in the Large Magellanic Cloud (LMC), it is approximately 23 6 (Dunstall et al., 2015).
@@
To address these problems, we study the multiplicity properties of massive BSGs ( ) in the Small Magellanic Cloud (SMC) within the new Binarity at LOw Metallicity (BLOeM, Shenar et al., 2024) survey. Apart from homogeneous and statistically significant numbers of observed massive stars, this survey gives the unique opportunity to study the evolution of BSGs at low metallicity (). The BLOeM survey contains multi-epoch spectroscopic observations of nearly 1000 stars in the SMC across different spectral types. The survey has been divided into different sub-samples that represent various evolutionary stages of massive stars across the HRD, namely: the O-type star domain (Sana et al., subm. in Nature Astronomy), the B-type dwarfs (spectral types B0-B3, see Villaseñor et al., 2025), OeBe stars (Bodensteiner et al., 2025), the BSGs (present work), and the cool supergiants (spectral types later than B3 up to F-type supergiants, see Patrick et al., 2025).
@@ 4.1 Multiplicity of BSGs
The overall determined spectroscopic binary fraction of BSGs in the SMC is similar to those detected in the LMC and in the Galaxy. Dunstall et al. (2015) analyzed multiplicity properties of 408 B-type stars in the LMC and reported the observed binary fraction among non-supergiants to be 25 2% and 23 6% for supergiants. In the Galaxy, de Burgos et al. (2025) found an average 27% of SB1 systems within a large statistical sample of B0-B6 supergiants. Such an independence of BSG binary fraction with respect to the metallicity is in agreement with the binary fraction of early B-type eclipsing systems in the Galaxy, LMC, and SMC (within the ¡ 20 d, see Moe & Di Stefano, 2013). As Moe et al. (2019) argued, the explanation for this could be that significant accretion rates in the massive star domain are so high that the massive protostellar disks will always become gravitationally unstable and fragment independent of the metallicity environment (which is not the case in the solar-type star domain). Furthermore, in the early B-type giants and dwarf BLOeMs’ sample, it has been found that binary fraction to some extent is dependent on metallicity (for details, see Villaseñor et al., 2025). In the next subsection, we discuss the binary fraction as a function of SpT and luminosity.
@@ 4.2 The BSGs landscape on the HRD
In addition, if we look at the distribution of spectroscopic binaries for a given spectral type (see the left panel in Fig. 10), we can notice that the B2 domain contains the largest number of spectroscopic binaries (especially ”lpv/SB1”) with respect to the rest of the spectral types in our BSG sample. In total, there are up to 50 of potential spectroscopic binaries (apart from the evolutionary effect), which can also be explained by their relatively low luminosity; thus, some of these targets could belong to the B-dwarf sample, where the observed multiplicity fraction is significant (50, see the accompanying BLOeM paper Villaseñor et al., 2025). Also, some of these least luminous BSGs could be Cephei pulsators, as these tend to be more commonly found among late-O and early-B spectral types (Burssens et al., 2020); thus, this could explain a given overpopulation of lpv/SB1 targets.
@@
Another reason for non-uniform distribution could be the selection bias of BSGs with the luminosity class of II, indeed, in this domain there is a significant overlap with the B-giants BLOeM sample (i.e., luminosity class of III, see Villaseñor et al., 2025), which could bias the overall number of BSGs in a given SpT and LC. Nevertheless, there is a significant drop in binary fraction below 18 kK and even though the present physical parameters of BSGs are not well defined yet, we can argue that it is a real drop in the binary fraction of BSGs at a given effective temperature and spectral type. Moreover, RV variations of late B-type supergiants are significantly smaller (less than 10 km s-1 for SpT B5 and later; see Patrick et al., 2025) than the ones studied in the present work, which confirms the absence of short-period supergiants at SpT later than B3. Nevertheless, the binary fraction among B3 BSGs is 20%, which agrees with the observed spectroscopic binary fraction of more evolved late-B, A-, and F-type supergiants (20% for the B5-9s SpT domain and significantly smaller binary fraction for the A- and F-type supergiants). It could indicate that at this SpT domain (B2) and regime, we observe the end of MS for a studied sample of BSGs; actually, it will also be interesting to see the behavior of sin for the same targets. If there are no fast-rotating ( sin ¡ 100 km s-1) BSGs with a SpT later than B2, it can also support the idea of the MS boundary at this evolutionary phase (as it has been shown in Vink et al., 2010). Notably, the BLOeM survey has been designed in a way to observe as many as possible bright massive stars in the SMC, while also aiming to achieve a homogeneous sampling across the G-band from 10 to 16.5 mag. The overall sample completeness for the stars with (i.e. those born as B-type stars) is 20% with respect to the Gaia catalogue within the range of 13 – 15 mag in the G-band (see Fig. 2 in Shenar et al., 2024). Thus, the given sample of BSGs and the other sub-samples of the BLOeM survey do not have a significant sample selection bias in terms of the magnitude and colour of targets that would affect the interpretation of our results because of the incompleteness of the samples.
@@ Acknowledgements.
- Villaseñor et al. (2025) Villaseñor, J. I., Sana, H., Mahy, L., et al. 2025, arXiv e-prints, arXiv:2503.21936